# Exploring Holomorphic Retracts

### v6

@@ 1. Introduction
Given this, it may seem questionable as to why we have stated theorem 1.3 in this introductory section; in this regard let us highlight the existence of domains as already alluded to, whose boundaries are not holomorphically extreme but are $\mathbb{C}$-extremal. Also, $\mathbb{C}$-extremality is far easier to verify than holomorphic extremality which apparently is a more restrictive condition. Indeed, a concrete example of a domain of the kind just mentioned is recorded in proposition 8.8 wherein the domain also serves to show that the converse of the above proposition is false. As for the significance of the above proposition 1.4 itself, let us only remark that it will be employed in determining retracts through the origin of all possible dimensions in $\ell_{q}$-‘balls’ as in theorem 1.9 below. It also helps in establishing biholomorphic inequivalences, particularly with product domains; for instance, a domain as in the above proposition with holomorphically extreme boundary and the polydisc – this is only to give a quick example here. To get new results, we must consider domains with greater complexity, to proceed systematically towards which we deal first with a simple class of domains which allow analytic varieties in their boundaries. Stating the essence here differently, we may use knowledge about retracts of a domain to detect whether it is irreducible i.e., biholomorphic to a product domain. Let us only indicate the key idea here namely, the fundamental common feature amongst all product domains that: through every point of a product domain, there are sufficiently many directions along which existence of non-trivial retracts is guaranteed. Let us just state precisely here the simplest result to this effect to indicate what this means; leaving more general formulations to be laid out in section 8 wherein we shall have other remarks to tie up with.
@@
Let $D_{1},D_{2}$ be domains in $\mathbb{C}$ with one of them bounded, say $D_{1}$. Then there exists $\epsilon>0$ such that: for every vector $v_{2}\in T_{p_{2}}D_{2}\simeq\mathbb{C}$ with $|v_{2}|<\epsilon$, there exists a retract $Z$ of $D_{1}\times D_{2}$ passing through $p=(p_{1},p_{2})$ with $(1,v_{2})\in T_{p}Z$.
@@
As already alluded to, this paves the way for establishing the following.
Proposition 1.6.
Let $D$ be any bounded balanced domain of holomorphy in $\mathbb{C}^{2}$ (whose Minkowski functional we denote by $h=h_{D}$), which is not convex at an open dense subset of $\partial D$ (that there exist plenty of domains will be discussed later); here by not ‘convex at a boundary point’ $p\in\partial D$, means that there are no ‘supporting linear functionals’ for $p$ i.e., there is no $\mathbb{C}$-linear functional $\ell:\mathbb{C}^{2}\to\mathbb{C}$ with $|\ell|\leq h$ and $|\ell(p)|=h(p)$. Let $A$ be any proper analytic subset of $D$ not passing through the origin. Then, $D\setminus A$ is not biholomorphic to a product domain.
Domains as in the above proposition may be regarded as being maximally non-convex in the realm of bounded balanced domains, remembering that it is impossible to have a bounded domain which is non-convex throughout the boundary i.e., at every boundary point. While convex domains is a much studied class, the other extreme has by far less been; this proposition contributes towards that. An extension of this to higher dimensions is briefly discussed in section 8. We also hasten to mention that there are better techniques to detect the occurrence of a product structure for domains far beyond the class of balanced domains owing particularly to the very recent work [Bh-Bo-Su], which however requires some quantitative knowledge about the squeezing function. However we believe that for balanced domains, the above proposition gives a much simpler approach in practice, since a balanced domain is most often given by its Minkowski functional. Needless to say, there are plenty of balanced domains of holomorphy satisfying the non-convexity condition in the above proposition, as will also be discussed in section 8.
@@
Getting back to retracts, let us start by noting that while Vesentini’s theorem says that all retracts of $D$ (as in that theorem) passing through the origin, are linear, it does not say that all linear subspaces are retracts. Infact, this is utterly false as soon as we consider any such $D$ as in the above theorem other than balls linearly equivalent to the standard Euclidean ball; this is due to the result of Bohnenblust in [Bhnblst] (prior to which the real version was obtained by Kakutani in [Kakutni]). The question of precisely which linear subspaces are realizable as retracts of $D$ as in the above theorem is significant for us. For instance: while every retract of the polydisc $\Delta^{N}$ is a graph, not every graph in $\Delta^{N}$ is a retract; only those which are graphs on linear retracts of $\Delta^{N}$ are, making it important to precisely pin down which linear subspaces are retracts. Indeed, this is quite an interesting question in its own right and is tied to a classical question in the geometry of Banach spaces which in our context of finite dimensional (complex) Euclidean spaces reads as: which are those linear subspaces of $\mathbb{C}^{N}$ equipped with some norm $\mu$, that are realizable as images of norm-one (linear) projections (when we measure the norm of the projection with respect to the operator norm induced by $\mu$)? The answer turns out to vary even among the $l^{p}$ norms i.e., depend upon $p$ when we take $\mu=\|\cdot\|_{l^{p}}$ even if we exclude the case $p=2$, which may not be surprising at all for this case is distinguished as the one that arises from an inner product; specifically, the case $p=\infty$ is also different from the other cases. Before getting to such differences, let’s mention a unified result. As our interest lies in retracts, we shall phrase in it in the language of retracts rather than norm-one projections. Specifically, the characterization of linear (thereby holomorphic, owing to Vesentini’s theorem) retracts of the standard $l^{p}$ ball ($1\leq p<\infty,~{}p\neq 2$) can be readily deduced from those of norm-one projections in $\mathbb{C}^{N}$ equipped with $l^{p}$-norm recorded in the well-known book on classical Banach spaces by Lindenstrauss – Tzafiri and it reads as follows for the finite dimensional case wherein the result holds for $p=1$ as well, unlike its infinite dimensional counterpart.
@@
$u_{j}=\sum_{i\in\sigma_{j}}\lambda_{i}e_{i},~{}1\leq j\leq m,~{}with~{}\sigma_{j}\cap\sigma_{k}=\phi~{}for~{}k\neq j$
@@
We shall discuss alternative formulations of the above result which are more geometrically revealing in section 5. In-particular, retracts of $\ell^{p}$-balls are isometric copies of lower dimensional standard $\ell^{p}$-balls. Next, we note that the case $p=2$ has been excluded only because in that case $B_{\ell^{2}}=\mathbb{B}$, in which all linear subspaces are retracts. We would rather like to draw attention to the more important fact that this result is not true for $p=\infty$ to demonstrate which (by a counterexample as simple as possible), we include an example in section 3. This naturally leads to the question as to whether it is possible at all to pin down precisely which linear subspaces of the polydisc $\Delta^{N}$ can be realized as linear retracts (recall that the polydisc is the exceptional case corresponding to $p=\infty$ which is left out by theorem 1.7). Towards this, we first note the following characterization of those linear subspaces of $\Delta^{N}$ which are its retracts. Indeed, this can readily derived from [Heath_Suff], though it is not stated in these explicit terms; contained within our treatment will be a simpler proof of the main result of [Heath_Suff] about the characterization of all retracts of the polydisc as well as a deduction of the following special one about its linear retracts.
@@
    For each $m\in\{1,2,\ldots,N\}\setminus J$, the scalars $c_{m,j}$, satisfy $\sum_{j\in J}|c_{m,j}|\leq 1$.
@@
It is immediate that through any given point of the polydisc, there exist retracts of all possible dimensions ($k=0,1,\ldots,N$), as its automorphism group acts transitively; a little contemplation shows that the transitive action is unnecessary to confirm the existence of such retracts. Indeed, for $1\leq p<\infty$, the $\ell_{p}$-ball fails to be homogeneous unlike the polydisc as soon as $p\neq 2$; but it is easy to derive as a corollary of theorem 1.7 that through each point of the $\ell_{p}$-ball, there exist retracts of all possible dimensions, as the retracts through the origin serve to confirm this. It is too much to as for retracts along any given set of directions because as the above proposition shows: not every two dimensional subspace is a retract, a fact which holds not only for $\Delta^{N}$ but also for every ball that is not linearly equivalent to the ball with respect to the $l^{2}$-norm, $\mathbb{B}^{N}$. We must not be amiss to note however that every one dimensional linear subspace of a bounded balanced convex domain in $\mathbb{C}^{N}$ is a retract – this follows by the Hahn–Banach theorem. It is natural to ask whether this holds if we relax the convexity condition on the domain to pseudoconvexity. The answer is in the negative, as was noted by Lempert in [Lempert-FundArt] by indicating an example to complement his passing observation therein that the existence of $2$-dimensional retracts is rather restricted and to the effect that when convexity is relaxed, even $1$-dimensional retracts can fail to exist in abundance. The only rigorous record of the just-mentioned Lempert’s example is in [Jrncki_invrnt_dst] (specifically in example 11.3.10 therein). While this relies on the usage of complex geodesics and is done only for one particular direction, we found that we can record a simpler example wherein we can begin to concretely address the question about retracts of higher dimension as well; moreover this can be achieved entirely through elementary means. In-fact, our example is as simple as the standard $\ell^{q}$-‘ball’ for $q<1$ and we record this in the following proposition; an interesting feature of the $\ell_{q}$-‘ball’ in $\mathbb{C}^{N}$ is that apart from failing to be convex, it fails to be concave unlike its real counter-part. Let us also note that while the above-mentioned example of Lempert is a pseudoconvex complete Reinhardt domain much as ours, we believe recording the aforementioned $\ell^{q}$-‘ball’ is worthwhile due to the fact that the literature about retracts of dimension higher than one, is scant at-least compared to complex geodesics and one-dimensional retracts. In that regard, we would like to mention the very recent article of [GZ] which appeared on the arXiv, a couple of months after the earlier versions of our article. Ghosh – Zwonek attain in this article [GZ], a complete characterization of the two dimensional retracts (and thereby all retracts) of the three dimensional Lie ball and the tetrablock; this article also has noteworthy general results about retracts of Lempert domains. We now get to the special case of the $\ell^{q}$-‘ball’ for $q<1$ in arbitrary (finite) dimension, which is amongst the simplest examples of (pseudoconvex) domains which fail to be a Lempert domain (see proposition 5.8 below as well); indeed, we deal with large collections of domains that fail to be Lempert domains (however, it goes without saying that the case of Lempert domains is indeed of interest).
@@
Let $q$ be an arbitrary positive number less than one. Consider the $\ell^{q}$-‘ball’ given by $D_{q}=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}\;:\;|z_{1}|^{q}+\ldots+|z_{N}|^{q}<1\}$ and the $r$-dimensional subspace $Y=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}\;:\;z_{i}=a_{1,i}z_{1}+\ldots+a_{r,i}z_{r}~{}\text{for}~{}r+1\leq i\leq N\}$ of $\mathbb{C}^{N}$, where for each fixed $i$, not all $a_{i,j}$’s are zero. Then the $r$-dimensional subspace $Y_{q}=D_{q}\cap Y$ is not a retract of $D_{q}$.
In-fact, the only $r$-dimensional linear retracts of $D_{q}$ are those spanned by an $r$-element subset of the standard basis $\{e_{1},\ldots,e_{N}\}\subset\mathbb{C}^{N}$. Thus, the holomorphic retracts through the origin of $D_{q}$ are precisely those linear subspaces spanned by some subset of the standard basis vectors.
@@
Note that the boundaries of these domains are holomorphically extreme; indeed $\mathbb{C}$-extremal, as is the case for $\ell^{p}$-balls for $1\leq p<\infty$ as well, in which case Vesentini’s theorem reduces considerations from holomorphic retracts to linear ones. We also furnish examples (of balanced non-convex pseudoconvex domains) for the case where such extremality condition fails in the section where we detail the proof of the above theorem. Indeed, we shall show in such examples that retracts of all possible dimension exists through each point of the domain other than the origin as well. That retracts cannot also be expected in any given direction for the non-convex case, is substantiated by $D_{q}$ as in the above theorem, owing to the spectacular failure even at the origin (of $1$-dimensional retracts for instance, along any direction other than the coordinate directions). Getting back to analysing a bit further, examples arising out of $\ell^{p}$-balls, let us remark a word or two, about concrete examples of domains satisfying the $\mathbb{C}$-extremality condition in his result before we pass out of it. Firstly, while the product operation fails to generate further examples of balls with $\mathbb{C}$-extremal property, intersection does i.e., the intersection of a pair of balls, possibly with respect to different norms, with $\mathbb{C}$-extremal boundary is again a ball (with respect to some norm) with $\mathbb{C}$-extremal boundary; Vesentini’s theorem applies to pin-down the retracts through the center of such a ‘ball of intersection’. Finally and more importantly, to bring out the role of the extremality condition both in Vesentini’s theorem and allied results both in the foregoing as well as the sequel (such as theorem 1.15 involving $\mathbb{R}$-extremality), let us mention that the converse of Vesentini’s theorem is also true: if all retracts through the origin of a bounded balanced convex domain are linear, then the domain is $\mathbb{C}$-extremal. In-fact, it is enough to just have linearity of the one-dimensional retracts and furthermore, a general and local version can be formulated as follows.
@@
    That the hypotheses that $\partial D$ is convex near $p$ implies convexity of $\partial D$ at $p$ together-with the continuity of the Minkowski functional near $p$; this will be discussed during the course of the proof. It may happen that $\partial D$ is convex at some boundary point but not in any open piece of the boundary about it; to see a simple example, consider $l^{q}$-‘balls’ for $q<1$ in $\mathbb{C}^{2}$ and examine points $p$ on the coordinate axes. This gives an example wherein convexity near $p$ fails but the weaker condition of convexity at $p$ together-with the continuity of the Minkowski functional near $p$ indeed holds. Finally here let us recall that if $p$ is a non-convex boundary point then there are no one-dimensional retracts through the origin in the direction of $p$.
@@
    More importantly, a precise description of all the non-linear retracts through the origin in a bounded balanced pseudoconvex domain along the direction of a boundary point $p$ at which $\partial D$ is convex but fails to be $\mathbb{C}$-extremal – with no assumptions about the continuity of the Minkowski functional near $p$ – is attained in proposition 6.4. We record the essence of this in the following theorem.

### v7

@@ 1. Introduction
We shall employ this in determining retracts through the origin of all possible dimensions in $\ell_{q}$-‘balls’ as in theorem 1.8 below. The $\ell_{q}$-balls give the simplest domains of holomorphy with holomorphically extreme but non-convex boundaries, at-least for the case where global $C^{1}$-smoothness of boundary fails. Now, the above proposition also helps in establishing biholomorphic inequivalences, particularly with product domains; for instance, a bounded balanced pseudoconvex domain with holomorphically extreme boundary, and the polydisc – this is only to give a quick example here. We now discuss other possible results about biholomorphic inequivalences, but only briefly as the focus of this article is more on attaining descriptions of retracts. Towards this, we consider domains with greater complexity to proceed systematically towards which, we deal with the simplest class of domains generalizing the polydisc other than products, that are also rich in analytic varieties in their boundaries, namely analytic polyhedra. Now, stating the essence of the foregoing differently, we may use knowledge about retracts of a domain to detect whether it is irreducible i.e., not being biholomorphic to a product domain. The key idea here is the the following common feature amongst all product domains: through every point of a product domain, there are sufficiently many directions along which existence of non-trivial retracts is guaranteed. Let us just state here the simplest result to this effect to indicate what this means; leaving more general formulations to be laid out in section 8 wherein we shall have other remarks to tie up with and furnish applications.
@@
Let $D_{1},D_{2}$ be domains in $\mathbb{C}$ with one of them bounded, say $D_{1}$. Let $p$ be an arbitrary point of $D=D_{1}\times D_{2}$. Then there exists $\epsilon>0$ such that: for every vector $v_{2}\in T_{p_{2}}D_{2}\simeq\mathbb{C}$ with $|v_{2}|<\epsilon$, there exists a retract $Z$ of the product domain $D$ passing through $p=(p_{1},p_{2})$ with $(1,v_{2})\in T_{p}Z$.
@@
The version of this for arbitrary (finite) dimensions is formulated in proposition 8.5 which as already alluded to, paves the way for establishing the following.
Theorem 1.5.
Let $D$ be a (bounded) balanced $\mathbb{C}$-extremal analytic polyhedron. Let $G$ be any domain in $\mathbb{C}^{N}$ admitting retracts for some open piece of directions at each $w\in G$. Then $D$ is not biholomorphic to $G$. The same is true when $D$ is replaced by $D\setminus A$ as well, where $A$ is any non-trivial analytic set (possibly of codimension one) not containing the origin. In particular, $D$ and $D\setminus A$ are irreducible i.e., not biholomorphic to a product domain.
The precise meaning of what it means for a domain $G$ to admit retracts through a point $w\in G$ for some open piece of directions (based at the origin in $T_{w}G$) is laid out in definition 5.9. The condition of $\mathbb{C}$-extremality means that there are no (non-trivial) complex line segments in the boundary. As the foregoing contains results involving the conditions of $\mathbb{C}$-extremality and holomorphic extremality, let us mention here that: while the two notions are equivalent in the presence of convexity, this is not so in general, as is manifestly seen in the class of analytic polyhedra. The inequivalence of analytic polyhedra with strictly pseudoconvex domains (one of the simplest classes of domains with holomorphically extreme boundaries) is well-known due to Henkin’s 1973 article [henkin]. Product domains also share the same property: none of them is biholomorphic to any domain with holomorphically extreme boundary; the techniques of Remmert – Stein attribute such inequivalences to the lack of analytic varieties in the boundaries of domains in the latter class. The question about the irreducibility of analytic polyhedra $D$ and associated ‘co-analytic domains’ $D\setminus A$ for analytic sets $A\subset D$, particularly when the codimension of $A$ is one, is therefore interesting. One cannot hope for results without imposing conditions as there are many analytic polyhedra which are completely reducible, the polydisc being the simplest example. The above result shows that irreducibility of $D$ is guaranteed, as soon as the milder $\mathbb{C}$-extremality condition holds at-least within the class of balanced domains. We hasten to mention that while there may well be shorter proofs than the one via retracts to prove the irreducibility of those analytic polyhedra $D$ as in the above theorem, it gets somewhat more difficult firstly for the associated co-analytic domains $D\setminus A$ therein; and all the more difficult for more general cases. As our purpose in this article is to attain results characterizing retracts more than various applications, we do not strive to include beyond a couple of results of this kind. We also hasten to mention that there are better techniques to detect the occurrence of a product structure for domains far beyond the class of balanced domains owing particularly to the very recent work [Bh-Bo-Su], which however requires some quantitative knowledge about the squeezing function. However, we believe that the approach via retracts is worth recording, as there are no universal techniques to establish biholomorphic inequivalences. We shall indeed show that there are plenty of balanced polyhedra as in the above theorem, in chapter 8.
@@
Getting back to retracts, let us note that while Vesentini’s theorem says that all retracts of $D$ as in that theorem, passing through the origin, are linear, it does not say that all linear subspaces are retracts. Infact, this is utterly false as soon as we consider any such $D$ as in the above theorem other than balls linearly equivalent to the standard Euclidean ball; this is due to the main result of Bohnenblust in [Bhnblst] (prior to which the real version was obtained by Kakutani in [Kakutni]). The question of precisely which linear subspaces are realizable as retracts of $D$ as in Vesentini’s theorem is significant for us. For instance: while every retract of the polydisc $\Delta^{N}$ is a graph, not every graph in $\Delta^{N}$ is a retract; only those which are graphs on linear retracts of $\Delta^{N}$ are, making it important to first precisely pin down which linear subspaces are retracts. Indeed, this is quite an interesting question in its own right and is tied to a classical question in the geometry of Banach spaces which in our context of finite dimensional (complex) Euclidean spaces reads as: which are those linear subspaces of $\mathbb{C}^{N}$ equipped with some norm $\mu$, that are realizable as images of norm-one (linear) projections when we measure the norm of the projection on $\mathbb{C}^{N}$ with respect to the operator norm induced by $\mu$? There are differences in the answers for the cases $p=1,\infty$ even if we exclude $p=2$, the well-known case distinguished as arising from an inner product, thereby not surprising at all. Before getting to such differences, let us mention that the result is uniform for all finite $p$ other than two and just state this unified (known) result. As our interest lies in retracts, we shall phrase it in the language of retracts rather than norm-one projections. Specifically, the characterization of linear retracts (thereby those holomorphic retracts which pass through the origin, owing to Vesentini’s theorem) of the standard $l^{p}$ ball denoted $B_{\ell^{p}}$, can be readily deduced from those of norm-one projections in $\mathbb{C}^{N}$ equipped with $l^{p}$-norm recorded in the well-known book on classical Banach spaces by Lindenstrauss – Tzafiri and it reads as follows for the finite dimensional case wherein the result holds for $p=1$ as well, unlike its infinite dimensional counterpart.
@@
(1.1) $u_{j}=\sum_{i\in\sigma_{j}}\lambda_{i}e_{i},~1\leq j\leq m,~with~\sigma_{j}\cap\sigma_{k}=\phi~for~k\neq j$
@@
We shall derive in section 5, alternative formulations (some of which are known) of the above result, which are more geometrically revealing to show that the linear retracts of $\ell^{p}$-balls are precisely isometric copies of lower dimensional standard $\ell^{p}$-balls. Stated differently, only those linear subspaces of $\ell^{p}$-balls which are biholomorphic to lower dimensional $\ell^{p}$-balls, turn out to be their holomorphic retracts through the origin. Next, we note that the case $p=2$ has been excluded only because in that case $B_{\ell^{2}}=\mathbb{B}$, wherein all linear subspaces are retracts. We would rather like to draw attention to the more important fact that this result is not true for $p=\infty$ to demonstrate which (by a counterexample as simple as possible), we include an example in section 3. Nevertheless, the aforementioned geometric characterization of the linear retracts of $\ell^{p}$-balls for finite $p$, as lower dimensional versions of themselves, holds true for the polydisc as well. Indeed, this can derived without much difficulty from [Heath_Suff], though it is not stated in these explicit terms.
@@
    For each $m\in\{1,2,\ldots,N\}\setminus J$, the scalars $c_{m,j}$ satisfy $\sum_{j\in J}|c_{m,j}|\leq 1$.
Consequently, non-trivial linear retracts of the polydisc $\Delta^{N}$ are precisely those linear subspaces which are isometric copies of lower dimensional polydiscs $\Delta^{k}$ (for $1\leq k\leq N$).
@@
It follows that all holomorphic retracts of $\Delta^{N}$ (not necessarily passing through the origin) are biholomorphic copies of lower dimensional polydiscs. It is immediate that through any given point of the polydisc, there exist retracts of all possible dimensions ($k=0,1,\ldots,N$), as its automorphism group acts transitively; a little contemplation shows that the transitive action is unnecessary to confirm the existence of such retracts. Indeed, for every $1\leq p<\infty$, the $\ell^{p}$-ball fails to be homogeneous unlike the polydisc as soon as $p\neq 2$; but it is easy to derive as a corollary of theorem 1.6 that through each point of the $\ell^{p}$-ball, there exist retracts of all possible dimensions, as the retracts through the origin serve to confirm this. It is too much to ask for retracts along any given set of directions because as the above proposition shows: not every two-dimensional subspace is a retract, a fact which holds not only for $\Delta^{N}$ but also for every ball that is not linearly equivalent to the ball with respect to the $l^{2}$-norm, $\mathbb{B}^{N}$. We must not be amiss to note however that every one dimensional linear subspace of a bounded balanced convex domain in $\mathbb{C}^{N}$ is a retract – this follows by the Hahn–Banach theorem. It is natural to ask whether this holds if we relax the convexity condition on the domain to pseudoconvexity. The answer is in the negative, as was noted by Lempert in [Lempert-FundArt] by indicating an example to complement his passing observation therein, that the existence of $2$-dimensional retracts is rather restricted i.e., to the effect that when convexity is relaxed, even $1$-dimensional retracts can fail to exist in abundance. The only rigorous record of the just-mentioned Lempert’s example is in [Jrncki_invrnt_dst] (specifically in example 11.3.10 therein). While this relies on the usage of complex geodesics and is done only for one particular direction, we found that we can record a simpler example wherein we can begin to concretely address the question about retracts of higher dimensions as well; moreover this can be achieved entirely through elementary means. In-fact, our example is as simple as the standard $\ell^{q}$-‘ball’ for $q<1$ and we record this in the following proposition; an interesting feature of the $\ell^{q}$-‘ball’ in $\mathbb{C}^{N}$ is that apart from failing to be convex, it fails to be concave unlike its real counter-part. Let us also note that while the above-mentioned example of Lempert is a pseudoconvex complete Reinhardt domain much as ours, we believe recording the aforementioned $\ell^{q}$-‘ball’ is worthwhile due to the fact that the literature about retracts of dimension higher than one, is scant at-least compared to complex geodesics and one-dimensional retracts. In that regard, we would like to mention the very recent article [GZ]  [^5] This appeared on the arXiv, a couple of months after the earlier versions of our article. Ghosh – Zwonek attain in this article [GZ], a complete characterization of the two dimensional retracts (and thereby all retracts) of the three dimensional Lie ball and the tetrablock; this article also has noteworthy general results about retracts of Lempert domains. We now get to the special case of the $\ell^{q}$-‘ball’ for $q<1$ in arbitrary (finite) dimension, which is amongst the simplest examples of (pseudoconvex) domains which fail to be a Lempert domain (see proposition 5.17 below as well); indeed, we deal with large collections of domains that fail to be Lempert domains.
@@
Let $q=(q_{1},\ldots,q_{N})$, where $q_{j}$’s are arbitrary positive numbers, all less than one. Consider the $\ell^{q}$-‘ball’ given by
(1.2) $D_{q}=\{(z_{1},\ldots,z_{N})\in\mathbb{C}^{N}\;:\;|z_{1}|^{q_{1}}+\ldots+|z_{N}|^{q_{N}}<1\}.$
The holomorphic retracts through the origin of $D_{q}$ are precisely the orthogonal projections of $D$ onto a span of a finite subset of the standard basis vectors of $\mathbb{C}^{N}$; thereby, the retracts are all copies of lower-dimensional $\ell^{q}$-balls.
@@
The techniques used to obtain this theorem can also be used to characterize retracts through the origin of some (bounded balanced pseudoconvex) non-decoupled domains as well (see corollary 5.11). Note that the boundaries of these domains are holomorphically extreme, in particular $\mathbb{C}$-extremal, as is the case for $\ell^{p}$-balls for $1\leq p<\infty$ as well; in such cases proposition 1.3 reduces considerations from holomorphic retracts to linear ones. We also include a couple of examples of certain balanced non-convex pseudoconvex domains to illustrate the case where extremality condition fails (in the same section 5 where we detail the proof of the above theorem). Indeed, we shall show in such examples that retracts of all possible dimension exists through each point of the domain other than the origin as well. That retracts cannot also be expected in any given direction for the non-convex case, is substantiated by $D_{q}$ as in the above theorem, owing to the spectacular failure even at the origin (of $1$-dimensional retracts, along any direction other than the coordinate directions).
Getting back to analysing a bit further, examples arising out of $\ell^{p}$-balls (where $p\geq 1$), let us remark a word or two, about generating examples of domains satisfying the $\mathbb{C}$-extremality condition in Vesentini’s result, before we pass out of it. Firstly, while the product operation fails to generate further examples of balls with $\mathbb{C}$-extremal property, intersection does i.e., the intersection of a pair of balls, possibly with respect to different norms, with $\mathbb{C}$-extremal boundary is again a ball (with respect to some norm) with $\mathbb{C}$-extremal boundary; Vesentini’s theorem applies to pin-down the retracts through the center of such a ‘ball of intersection’. Now towards non-convex cases, Vesentini’s theorem cannot in general be modified easily so as to adapt to other classes of domains despite proposition 1.3; for instance, to deal with those analytic polyhedra whose boundaries are $\mathbb{C}$-extremal. Such domains can never be holomorphically extreme (so proposition 1.3 cannot be applied), as their boundaries are Levi flat on an open dense subset  [^6] Indeed, this is the maximum possible Levi flatness that can be expected in this context; for instance, while it is possible for a smoothly bounded domain $D$ to be holomorphically extreme all along $\partial D$, it is impossible for $\partial D$ to be Levi flat everywhere.. However, we were able to attain the following result.
Theorem 1.9.
There do not exist any (non-trivial) retracts through the origin of (bounded) balanced $\mathbb{C}$-extremal analytic polyhedron $D$ in $\mathbb{C}^{N}$ along the direction of any of the points in the open faces of $\partial D$. To be more precise, suppose that $D$ can be described as
$D=\{z\in U\;:\;|f_{k}(z)|<1,\text{ for all }1\leq k\leq m\},$
for some holomorphic $f_{k}$’s in a neighbourhood $U$ of $\overline{D}$ (and that $D$ happens to be an open, connected, bounded, balanced subset of $\mathbb{C}^{N}$ with $\mathbb{C}$-extremal boundary). Suppose also that the system of defining functions $(f_{k})$ is minimal i.e., none of the functions $f_{k}$ can be removed without changing $D$. Let $p$ be a boundary point on any of the open faces of $\partial D$, i.e. $p\in\tilde{\sigma}_{j}$ for some $j$ with $1\leq j\leq m$, where
$\tilde{\sigma}_{j}:=\{z\in\partial D:|f_{j}(z)|=1,~|f_{l}(z)|<1~\text{for}~1\leq l\leq m~\text{with}~l\neq j\}.$
Then there does not exist any retract $Z$ of $D$ passing through the origin tangent to $p$ thereat i.e., with $p\in T_{0}Z$.
We know that $T_{0}Z$ is itself a retract, indeed a linear retract. What the above theorem means for linear retracts of $D$ is that they cannot intersect the open faces $\tilde{\sigma}_{j}$. Thereby, every (non-trivial) linear retract is spanned by (at-most N1N-1) points from the lower dimensional ribs of the polyhedron. Consequently, owing to theorem 1.15 below, every (non-trivial holomorphic) retract through the origin is a biholomorphic copy of a lower dimensional linear slice of the same polyhedron spanned by vectors from its ribs.
Existence of domains as in the above theorem in abundance will be shown in section 8. A complete determination of retracts (even through the origin) of such domains seems possible only when either $f_{j}$’s are explicitly given or when sufficient data about them is given. We shall do this exercise only for what is perhaps the simplest of such polyhedrons namely, $D_{h}:=\{(z,w)\in\mathbb{C}^{2}:|z^{2}-w^{2}|<1,\;|zw|<2\}$. It may be noted that this itself is interesting to start with, as its boundary is not entirely complex non-degenerate, (i.e. its ribs has singular points within it).
The next main result of this paper, from which we deduce both the above theorem as well as theorem 1.8, is the following.
Theorem 1.10.
Let $D$ be a bounded balanced pseudoconvex domain in $\mathbb{C}^{N}$ and $C$ be the convex hull of $\overline{D}$. Let $L$ be any linear subspace which cannot be precisely spanned by vectors from $\partial D\cap\partial C$. Then $D_{L}:=D\cap L$ is not a linear retract of $D$.
We now lay down the aforementioned application of theorem 1.8 and theorem 1.9 about retracts of $D$ to establish irreducibility, i.e., biholomorphic inequivalence with product domains. As a consequence of the foregoing pair of theorems, we see that a holomorphic retract $Z$ through the origin of an analytic polyhedron as in the above theorem 1.9 has the property that $T_{0}Z$ intersects $\partial D$ only in its closed nowhere dense subset comprising ribs (formed by intersections of at-least two of the faces) of $\partial D$. Thereby, such retracts are sparse: $D$ does not admit any retract along any open piece of directions at the origin. This leads to theorem 1.5; though this was already stated, the point here is to highlight that the proof is very similar for theorem 1.8 we well. Indeed, to minimize repetition, we prove both cases simultaneously in section 8.
Corollary 1.11.
Let $D$ be a non-convex decoupled egg domain $D$ as in (1.2) of theorem 1.8. Then $D$ is not biholomorphic to a product domain. If $A$ is a non-trivial analytic subset of $D$ not passing through the origin then the same conclusion holds for $D\setminus A$.
Among others, the above corollary alongwith theorem 1.5 deals with three classes of domains: product domains, analytic polyhedra and ‘co-analytic’ domains  [^7] This is not standard terminology; however, this helps prevent repeating long phrases to describe a class of domains that will occur oftentimes in this article. We shall also refer to them as ‘analytic complements’ at times. i.e., domains of the form $D\setminus A$ wherein $A$ is a non-trivial analytic subset of $D$ (we only deal with domains of holomorphy $D$ assuming ${\dim A}>0$ needless to say). Members of all these classes, share the common feature that their boundaries have non-trivial analytic varieties in them. Corollary 1.11 is to the effect that nevertheless it is easy to get hold of mutually inequivalent members from these classes. In-fact, as we shall see along the way in section 8 (one of the major sections of this article), an analytic polyhedron is never equivalent to a co-analytic domain $D\setminus A$, whatever $D$ may be. As can be guessed, such inequivalences stem from the differences in the location of the analytic boundary pieces (for the former, these are not situated in the interior as is evidently the case with the latter). An analytic polyhedron can be equivalent to a product domain, the simplest example being $\Delta^{2}$; however, proposition 8.11 will confirm that this is the only bounded balanced domain in $\mathbb{C}^{2}$ for which such an equivalence occurs. Finally and more importantly, to bring out the role of the extremality condition both in Vesentini’s theorem and allied results (such as the foregoing theorem 1.10 and theorem 1.17 below), let us mention that the converse of Vesentini’s theorem is also true: if all retracts through the origin of a bounded balanced convex domain are linear, then the domain is $\mathbb{C}$-extremal. In-fact, it is enough to just have linearity of the one-dimensional retracts and furthermore, a general and local version can be formulated, namely as follows.
@@
    We note that the hypotheses that $\partial D$ is convex near $p$ implies convexity of $\partial D$ at $p$ together-with the continuity of the Minkowski functional near $p$; this will be discussed during the course of the proof. When $\partial D$ is convex near $p$, it is enough to check for the existence of some holomorphic variety through $p$ rather than a linear one, in order to ascertain existence of non-linear retracts; this is due to the equivalence of $\mathbb{C}$-extremality with holomorphic extremality in the setting of the above theorem, which is discussed in proposition 2.37. It may happen that $\partial D$ is convex at some boundary point but not in any open piece of the boundary about it; to mention a simple example, we may consider $l^{q}$-‘balls’ for $q<1$ in $\mathbb{C}^{2}$ and examine points $p$ on the coordinate axes. This gives an example wherein convexity near $p$ fails but the weaker condition of convexity at $p$ together-with the continuity of the Minkowski functional near $p$ as in the statement of the above theorem, indeed holds. Finally here let us recall that if $p$ is a non-convex boundary point then there are no one-dimensional retracts through the origin in the direction of $p$.
@@
    More importantly, a precise description of all the one-dimensional non-linear retracts through the origin in a bounded balanced pseudoconvex domain along the direction of a boundary point $p$ at which $\partial D$ is convex but fails to be $\mathbb{C}$-extremal – with no assumptions about the continuity of the Minkowski functional near $p$ as in the statement of the above theorem – is attained in proposition 6.4. That leads to the following theorem about retracts of all dimensions.